# Simulating the Visual World with Artificial Intelligence: A Roadmap

### v2

@@ 4.2 Generation 2 - Interactiveness: Controllability and Interactive Dynamics
Compared with Generation 1, Generation 2 world models achieve a marked leap in interactiveness, representing a decisive step toward dynamic and flexible interactive world modeling. Within this generation, interaction can be realized either through spatial conditions (e.g., sketches, depth, motion pose) or through navigation modes (e.g., trajectory, action, instruction). Spatial conditioned approaches [sketchvideo, das, bian2025gsdit, li2025pisa, xue2025phyt2v, wang2025wisa] excel in producing high‑quality videos with strong 3D dynamics, accurate condition–video consistency, and stable scene interactions, though their flexibility in real‑time control remains limited; nonetheless, these strengths firmly place them within the Generation 2 category. In addition, navigational world models [xiang2024pandora, aether, hong2024slowfast, hou2024camtrol, ren2025gen3c, he2025cameractrl2, fu20243dtrajmaster] offer greater flexibility and transferability, as they do not depend on pre‑existing contextual knowledge and can generalize across diverse entities and scenarios, making them inherently suitable for real‑time and general‑purpose applications.
@@ 4.2.1 Foundation Models with Dynamic-Aware World Knowledge
Building upon the first generation, Generation 2 foundation models make a decisive leap in interactiveness, representing a major step toward more dynamic and flexible world modeling. These advances stem from large‑scale pretraining, higher‑capacity architectures, and richer conditioning interfaces, enabling the models to adapt more effectively to downstream control mechanisms. As a result, spatial conditions and navigation‑mode controls can now be integrated with higher video fidelity, stronger spatiotemporal consistency, text-video consistency, greater control flexibility, and better 3D and semantic-level interaction, qualities that directly improve performance in downstream tasks across diverse scenarios.
@@ 4.2.4 Video Generation as World Model in Autonomous Driving
Autonomous driving represents another critical application domain for world models. Existing world models in this context are primarily employed to synthesize large-scale, multi-scene datasets or to perform visual and policy prediction tasks, serving as potential solutions to advance autonomous driving toward Level 4 autonomy.
@@
However, because such layout information often relies on annotations available only in curated datasets, these approaches may lack practicality for real-time deployment. Accurate and timely collection of layout annotations in real-world applications remains challenging. Therefore, we consider layout-based methods to be an early-stage bridge toward building full-fledged world models in the autonomous driving domain.
@@
Hybrid Navigation World Model. With the rapid advancement of autonomous driving, there is a growing trend toward the development of unified navigation world models in this domain. Recent works [russell2025gaia2, ni2025maskgwm, gao2024vista] have begun to support multiple conditioning signals and navigation modes simultaneously, e.g., [russell2025gaia2] combining text instructions with action signals, text with trajectories, or [ni2025maskgwm, gao2024vista] even integrating all four modalities: text instructions, trajectories, action commands, and goal point images. This convergence indicates that autonomous driving world models are gradually evolving toward planning capabilities, aligning with the characteristics of third-generation world models as defined in this paper.
@@ 4.2.5 Video Generation as World Model in Gaming
Controller Navigation World Model. In the gaming domain, key control represents a crucial navigation mode for world models. This includes both inputs from game controllers [yang2024playgen, bruce2024genie, kanervisto2025wham, valevski2024gamengen, chen2025maag, li2025hunyuangamecraft] and keyboard-based [che2024gamegenx, matrix, guo2025mineworld, yu2025gamefactory, xiao2025worldmem, alonso2024diamond, parker2024genie2, decart2024oasis] interactions. For instance, models [chen2025maag, valevski2024gamengen, kanervisto2025wham] rely on controller inputs for gameplay interaction, while GameGen-X [che2024gamegenx] utilizes keyboard keystrokes to achieve video-level control. Other models [guo2025mineworld, yu2025gamefactory, xiao2025worldmem, alonso2024diamond, decart2024oasis] adopt a combination of keyboard and mouse inputs to enable frame-level control over game environments. Naturally, the complexity and granularity of supported actions vary across these models.
Notably, Google’s Genie series [bruce2024genie, parker2024genie2] marks a significant advancement in this area. Genie [bruce2024genie] is the first generative interactive environment trained in an unsupervised manner using unlabelled Internet videos. It can synthesize videos frame by frame from diverse inputs, including text-to-image prompts, hand-drawn sketches, text descriptions, or even real-world photos, all of which can be modulated via real-time action inputs. Its successor, Genie 2 [parker2024genie2], serves as a large-scale foundation world model. It demonstrates emergent capabilities at scale, such as object manipulation, complex character animation, physical reasoning, and even the prediction of other agents’ behavior. The Genie series is thus progressively evolving toward the vision of an interactive open-world simulation model that underpins next-generation world modeling.
@@ 4.3 Generation 3 - Planning: Modeling the Future Evolution of Complex Systems
In the third generation of world models [nwm, assran2025vjepa2, shin2025motionstream, zhu2025astra], planning emerges as the core capability that begins to bridge the gap between environment-centered physical simulation and agent-centered reasoning. This stage encompasses both objective physical planning, such as navigation and scene evolution based on global environmental dynamics, and the potential of subjective or agentic planning, where the model internally simulates trajectories, intentions, or counterfactual futures from a first-person perspective. In this sense, planning represents the first step toward integrating a mental world model component within a primarily physical framework, enabling the model to reason not only about how the world evolves, but also about why such evolutions may occur under different beliefs or goals.
@@
Only a limited number of methods currently reach this level of capability. The boundary between Generation 2 and Generation 3 is marked by the emergence of real-time inference ($\geq$ 24 FPS) and navigation-aware generation, where the system can adaptively simulate world trajectories based on external or internal control signals. Models that demonstrate temporally consistent control and near-real-time rendering are regarded as exhibiting emergent potential toward this generation, even if fully consistent physics and complex multi-step planning have not yet been achieved.
@@
Despite these advances, current models remain limited in scalability, action richness, and long-duration temporal consistency. Nevertheless, planning in Generation 3 signifies the critical turning point where world models evolve from purely physical simulators toward agentic and mental world modeling, unifying the objective and subjective dimensions of intelligent simulation.
@@ 4.4 Generation 4 - Stochasticity: Modeling Outlier and Low-Probability Events
While these efforts largely operate within the realm of human-scale physics, i.e., mesoscopic spatiotemporal modeling, we envision the advanced stage of Generation 4 models to extend planning capabilities toward both macroscopic and microscopic scales.
@@ 4.5 Insight: Beyond Generation
In addition, beyond purely visual simulation, the recent emergence of multimodal generation models [Veo325, pika23, imaginegrok, sora, meta2025vibes, runway2024gen3, wan22, luma2025ray3] with both vision and audio modalities indicates a clear trajectory toward unified audiovisual world models, where sound, narration, and perception are jointly simulated with physics-based consistency.
@@ 5.1 Two Complementary Development Directions
World Models for Dicision and Control. In parallel with the vision and simulation centric trajectory of world modeling, another line of research has evolved within the reinforcement learning and robotics communities, where the focus lies on using learned world dynamics to support decision-making and control. Rather than striving solely for pixel-level realism, these models emphasize predictive internal representations that enable agents to plan, imagine, and act within latent space. Through learning compact transition dynamics, they allow an embodied system to anticipate outcomes of hypothetical actions, perform long-horizon reasoning, and optimize policies before real-world execution. Notable examples such as PlaNet [hafner2019planet], Dreamer Series [hafner2019dreamer, hafner2020dreamer2, hafner2023dreamerv3], and related latent-dynamics models have demonstrated that predictive simulation and planning can effectively bridge perception and control, providing a complementary perspective to video-generation-based world modeling. Together, these directions illustrate how world modeling can advance along both simulation fidelity and decision-oriented reasoning, converging toward a more unified understanding of environment and agency.

### v3

@@ 4.2 Generation 2 - Interactiveness: Controllability and Interactive Dynamics
Compared with Generation 1, Generation 2 world models achieve a marked leap in interactiveness, representing a decisive step towards dynamic and flexible interactive world modeling. Within this generation, interaction can be realized either through spatial conditions (e.g., sketches, depth, motion pose) or through navigation modes (e.g., trajectory, action, instruction). Spatial conditioned approaches [sketchvideo, das, bian2025gsdit, li2025pisa, xue2025phyt2v, wang2025wisa] excel in producing high‑quality videos with strong 3D dynamics, accurate condition–video consistency, and stable scene interactions, though their flexibility in real‑time control remains limited; nonetheless, these strengths firmly place them within the Generation 2 category. In addition, navigational world models [xiang2024pandora, aether, hong2024slowfast, hou2024camtrol, ren2025gen3c, he2025cameractrl2, fu20243dtrajmaster] offer greater flexibility and transferability, as they do not depend on pre‑existing contextual knowledge and can generalize across diverse entities and scenarios, making them inherently suitable for real‑time and general‑purpose applications.
@@ 4.2.1 Foundation Models with Dynamic-Aware World Knowledge
Building upon the first generation, Generation 2 foundation models make a decisive leap in interactiveness, representing a major step towards more dynamic and flexible world modeling. These advances stem from large‑scale pretraining, higher‑capacity architectures, and richer conditioning interfaces, enabling the models to adapt more effectively to downstream control mechanisms. As a result, spatial conditions and navigation‑mode controls can now be integrated with higher video fidelity, stronger spatiotemporal consistency, text-video consistency, greater control flexibility, and better 3D and semantic-level interaction, qualities that directly improve performance in downstream tasks across diverse scenarios.
@@ 4.2.4 Video Generation as World Model in Autonomous Driving
Autonomous driving represents another critical application domain for world models. Existing world models in this context are primarily employed to synthesize large-scale, multi-scene datasets or to perform visual and policy prediction tasks, serving as potential solutions to advance autonomous driving towards Level 4 autonomy.
@@
However, because such layout information often relies on annotations available only in curated datasets, these approaches may lack practicality for real-time deployment. Accurate and timely collection of layout annotations in real-world applications remains challenging. Therefore, we consider layout-based methods to be an early-stage bridge towards building full-fledged world models in the autonomous driving domain.
@@
Hybrid Navigation World Model. With the rapid advancement of autonomous driving, there is a growing trend towards the development of unified navigation world models in this domain. Recent works [russell2025gaia2, ni2025maskgwm, gao2024vista] have begun to support multiple conditioning signals and navigation modes simultaneously, e.g., [russell2025gaia2] combining text instructions with action signals, text with trajectories, or [ni2025maskgwm, gao2024vista] even integrating all four modalities: text instructions, trajectories, action commands, and goal point images. This convergence indicates that autonomous driving world models are gradually evolving towards planning capabilities, aligning with the characteristics of third-generation world models as defined in this paper.
@@ 4.2.5 Video Generation as World Model in Gaming
Controller Navigation World Model. In the gaming domain, key control represents a crucial navigation mode for world models. This includes both inputs from game controllers [bruce2024genie, kanervisto2025wham, valevski2024gamengen] and keyboard-based [yu2025gamefactory, alonso2024diamond, parker2024genie2, decart2024oasis] interactions. For instance, models [valevski2024gamengen, kanervisto2025wham] rely on controller inputs for gameplay interaction, while other models [yu2025gamefactory, alonso2024diamond, decart2024oasis] adopt a combination of keyboard and mouse inputs to enable frame-level control over game environments. Naturally, the complexity and granularity of supported actions vary across these models.
Notably, Google’s Genie series [bruce2024genie, parker2024genie2] marks a significant advancement in this area. Genie [bruce2024genie] is the first generative interactive environment trained in an unsupervised manner using unlabelled Internet videos. It can synthesize videos frame by frame from diverse inputs, including text-to-image prompts, hand-drawn sketches, text descriptions, or even real-world photos, all of which can be modulated via real-time action inputs. Its successor, Genie 2 [parker2024genie2], serves as a large-scale foundation world model. It demonstrates emergent capabilities at scale, such as object manipulation, complex character animation, physical reasoning, and even the prediction of other agents’ behavior. The Genie series is thus progressively evolving towards the vision of an interactive open-world simulation model that underpins next-generation world modeling.
@@ 4.3 Generation 3 - Planning: Modeling the Future Evolution of Complex Systems
In the third generation of world models [nwm, assran2025vjepa2, shin2025motionstream], planning emerges as the core capability that begins to bridge the gap between environment-centered physical simulation and agent-centered reasoning. This stage encompasses both objective physical planning, such as navigation and scene evolution based on global environmental dynamics, and the potential of subjective or agentic planning, where the model internally simulates trajectories, intentions, or counterfactual futures from a first-person perspective. In this sense, planning represents the first step towards integrating a mental world model component within a primarily physical framework, enabling the model to reason not only about how the world evolves, but also about why such evolutions may occur under different beliefs or goals.
@@
Only a limited number of methods currently reach this level of capability. The boundary between Generation 2 and Generation 3 is marked by the emergence of real-time inference ($\geq$ 24 FPS) and navigation-aware generation, where the system can adaptively simulate world trajectories based on external or internal control signals. Models that demonstrate temporally consistent control and near-real-time rendering are regarded as exhibiting emergent potential towards this generation, even if fully consistent physics and complex multi-step planning have not yet been achieved.
@@
Despite these advances, current models remain limited in scalability, action richness, and long-duration temporal consistency. Nevertheless, planning in Generation 3 signifies the critical turning point where world models evolve from purely physical simulators towards agentic and mental world modeling, unifying the objective and subjective dimensions of intelligent simulation.
@@ 4.4 Generation 4 - Stochasticity: Modeling Outlier and Low-Probability Events
While these efforts largely operate within the realm of human-scale physics, i.e., mesoscopic spatiotemporal modeling, we envision the advanced stage of Generation 4 models to extend planning capabilities towards both macroscopic and microscopic scales.
@@ 4.5 Insight: Beyond Generation
In addition, beyond purely visual simulation, the recent emergence of multimodal generation models [Veo325, pika23, imaginegrok, sora, meta2025vibes, runway2024gen3, wan22, luma2025ray3] with both vision and audio modalities indicates a clear trajectory towards unified audiovisual world models, where sound, narration, and perception are jointly simulated with physics-based consistency.
@@ 5.1 Two Complementary Development Directions
World Models for Dicision and Control. In parallel with the vision and simulation centric trajectory of world modeling, another line of research has evolved within the reinforcement learning and robotics communities, where the focus lies on using learned world dynamics to support decision-making and control. Rather than striving solely for pixel-level realism, these models emphasize predictive internal representations that enable agents to plan, imagine, and act within latent space. Through learning compact transition dynamics, they allow an embodied system to anticipate outcomes of hypothetical actions, perform long-horizon reasoning, and optimize policies before real-world execution. Notable examples such as PlaNet [hafner2019planet], Dreamer Series [hafner2019dreamer, hafner2020dreamer2, hafner2023dreamerv3], and related latent-dynamics models have demonstrated that predictive simulation and planning can effectively bridge perception and control, providing a complementary perspective to video-generation-based world modeling. Together, these directions illustrate how world modeling can advance along both simulation fidelity and decision-oriented reasoning, converging towardsa more unified understanding of environment and agency.